OpenStudy (anonymous):

Will fan and medal! An auction website charges $1 for a bid. The bidding starts at 1¢ and goes up 1¢ at a time. A television that is worth $2000 is won, on average, with a bid of $160. You make one bid at random. Find the expected value of the outcome of the bid. (Write as an exact decimal, with a negative sign, if necessary.) Expected Value: $

OpenStudy (anonymous):

160 in 1¢ steps means 16,000 bids so you invest $1 with a 1/16000 probability of a $2000 return the expected value (in dollars) is ___ -1 + [(1/16000) * 2000)] this is why auction websites exist

OpenStudy (biohazard9064):

thats correct but I must say it looks a lot like the yahoo answer for the same question

OpenStudy (anonymous):

yeah thats the exact same answer from yahoo answers haha @BioHazard9064. but so what do i put in the answer box?

OpenStudy (biohazard9064):

So its a fill the blank question fun

OpenStudy (biohazard9064):

wow

OpenStudy (biohazard9064):

so if its -1+ [(1/16000) * 2000)] = -0.875 but that don't help much

OpenStudy (biohazard9064):

to be honest I cant solve this but I know someone who might

OpenStudy (biohazard9064):

@DaBest21 Need Some help

OpenStudy (anonymous):

ok

OpenStudy (biohazard9064):

Im stuck

OpenStudy (anonymous):

@em2000 does that alot

OpenStudy (anonymous):

it is a 1/16,000 chance of earning 1,999 while there is a 15,999 chance of losing a dollar so find 1/16000 which is 0.0000625 so do 1-0.0000625=.9999375 so .9999375*-1=-.9999375+(1999*.0000625=.1249375+-.9999375=0.875 @BioHazard9064 @madison.bush